A boat is traveling in a river at a speed of 9 miles per hour. The river has a current of 7 miles per hour. If the boat's destination is a dock located 5 miles upstream, what will be the total trip time for the boat?

Answer by ankor@dixie-net.com(22740)

A boat is traveling in a river at a speed of 9 miles per hour.
The river has a current of 7 miles per hour.
If the boat's destination is a dock located 5 miles upstream, what will be the total trip time for the boat?
:
The effective speed of the boat will be 9-7 = 2 mph
time = 5/2 = 2.5 hrs